OpenStudy (anonymous):

Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2. I've tried to do this over and over and I just can get it.

OpenStudy (anonymous):

\[2(-4)^2 - 2(2)^4 + 3^2 - (-4)^2 + 2^4\]all the terms with 16 in it cancel out

OpenStudy (anonymous):

\[=2*16 - 2*16 + 9 - 16 + 16 = 9\]

OpenStudy (anonymous):

Okay so the answer is 9? if so can you explain to me why?

OpenStudy (anonymous):

all the terms with 16 are cancelled. 2*16 - 2*16 = 0 and the last 2 -16 + 16 = 0

OpenStudy (anonymous):

equivalent to x^2 - z^4 + y^2 and x^2 = z^4

OpenStudy (anonymous):

Which make 9?

OpenStudy (anonymous):

all thats left over is y^2 = 9

OpenStudy (anonymous):

OHHH Okay :)